OpenStudy (anonymous):

In a right triangle ABC, the legs are AB = 5 and BC = 12, and in the right triangle DEF, the hypotenuse is DE = 10, and EF = 8. Which is DF + AC? 19 15 13 12

OpenStudy (anonymous):

19

OpenStudy (anonymous):

use pythagorean to find for ABC x^2 + 64 = 100, AC = sq rt 36 = 6 for DEF 144 + 25 = 169 = DF sq rt 169 = 13. 13 + 6 = 19

OpenStudy (anonymous):

thank yall :)!

OpenStudy (anonymous):

ac = sqrt(ab^2+bc^2) =13 df= sqrt(de^2-ef^2) = 6 ac+df = 19